OpenStudy (anonymous):

Air pressure may be represented as a function of height (in meters) above the surface of the Earth, as shown below. P(h)=Po*e^-0.00012h In this function, Po is the air pressure at the surface of the Earth, and h is the height above the surface of the Earth, measured in meters. At what height will the air pressure equal 65% of the air pressure at the surface of the Earth? A. .65m B. 3587.4m C. 3589.9m D. 5416.7m

OpenStudy (anonymous):

@ranga do you think you could help me with this problem please?

OpenStudy (anonymous):

0.65=e^-0.00012h ln(0.65)=ln(e^-0.00012h) ln(0.65)=-0.00012h h=ln(0.65)/-0.00012

OpenStudy (anonymous):

ln(e^x)=x

OpenStudy (anonymous):

I'm sorry @dhouser , but how exactly would I solve that?

OpenStudy (anonymous):

Would I just divide .65 by -.00012?

OpenStudy (shamil98):

ln means the natural log take the natural of (0.65) first. then divide it

OpenStudy (shamil98):

Punching that into your calculator you get 3589.86 rounding to the tenth 3589.9

OpenStudy (anonymous):

wouldn't it be negative because I divide ln(.65) by -.00012? thats what Dhouser told me earlier so I'm not sure :o

OpenStudy (shamil98):

i'm not sure the exact mechanics but it turns out to be positive

OpenStudy (ranga):

ln(.65) will be negative.

OpenStudy (anonymous):

oh but ln(.65) divided by -.00012 would be positive 3589.9 like shamil said, correct?

OpenStudy (ranga):

yes.

OpenStudy (shamil98):

yeah negative/negative = positive

OpenStudy (anonymous):

woohoo! thanks y'all :)